ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 34 Sadyrbaev, Felix, University of Latvia, Latvia Nonlinear separation of zeros Two nontrivial solutions x(t) and y(t) of a nonlinear second order ordinary differential equation are considered provided that the zero is a solution also. The estimation of the number of zeros of x(t) between two consecutive zeros of y(t) is made on a basis of the comparison of local properties of x(t) and y(t), expressed in terms of the respective linear equations of variations. The classical Euler theorem on separation of zeros of solutions of linear equations is included.